OpenStudy (anonymous):

when -3 is multiplied by the sum of a number and 8, the result is -2. find the number

OpenStudy (anonymous):

Let the number be x . Then the sum of number and 8 = (x+8) Now -3(x+8) = -2. hence solve for x.

OpenStudy (anonymous):

your amazing<3

OpenStudy (anonymous):

-3(x+8) = -2 3(x+8) = 2 3x = 2-24= -22 x=-22/3 I hope